# Quantum entanglement in a pure state of strongly correlated quantum impurity systems

### v2

@@ Abstract
We consider quantum entanglement in strongly correlated quantum impurity systems for states manifesting interesting properties such as multi-level Kondo effect and dual nature between itineracy and localization etc.. For this purpose, we set up a system consisting of one or two quantum impurities arbitrarily selected from the system as a subsystem, and investigate quantum entanglement with its environmental system. We reduce the pure state of interest as described above to the subsystem, and formulate quantum informative quantities such as entanglement entropy, mutual information and relative entropy. We apply them to quantum impurity systems consisting of several sites and the single impurity Anderson model to study the relationship between their states and the behaviors of quantum informative quantities. The obtained results suggest that the method proposed here is promising for elucidating the quantum entanglement of pure states in various quantum impurity systems.
@@ I Introduction
Therefore, in this paper, quantum impurities are modeled by Anderson impurities, and in section II, we consider the density operator describing the pure state of interest reduced to the subsystem consisting of one or two selected Anderson impurities in the system. We then formulate quantum informative quantities such as entanglement entropy, mutual information, and relative entropy. In sections III and IV, we apply the formulations in the previous section to quantum impurity systems consisting of several sites and the Single Impurity Anderson Model(SIAM) to study the relationship between their states and the behaviors of quantum informative quantities. The discussion and conclusions are provided in the last section V.
@@ II.2 Subsystem $A$ consisting of two selected quantum impurities
III Results of application to isolated multi quantum impurity systems
III.1 Isolated two quantum impurity system
@@
III.2 Isolated Hubbard chain
@@
The set of the quantum numbers $({\mathcal{S}}_{z},{\mathcal{S}},{\mathcal{Q}})$ of the ground state $\psi^{({\rm G})}_{{\mathcal{S}}_{z},{\mathcal{S}},{\mathcal{Q}}}$ is $({\mathcal{S}}_{z},{\mathcal{S}},{\mathcal{Q}})=\left(0,0,0\right)$ if $M$ is even and $({\mathcal{S}}_{z},{\mathcal{S}},{\mathcal{Q}})=\left(\pm\frac{1}{2},\frac{1}{2},0\right)$ if $M$ is odd.
@@ IV.2 Quantum entanglement in energy eigenstates depending on $N_{\rm nrg}$
We next consider two states with ${\mathcal{S}}_{z}={\mathcal{S}}$ and ${\mathcal{S}}_{z}=-{\mathcal{S}}$ for the lowest energy state $\psi^{(N_{\rm nrg},\rm G)}_{\mathcal{Q}}({\mathcal{S}},{\mathcal{S}}_{z})$ within ${\mathcal{S}}\neq 0$ and the corresponding appropriate ${\mathcal{Q}}$ for each $N_{\rm nrg}$, and then reduce these two states to the Anderson impurity system and consider the relative entropy $D\left(\rho^{(A)}_{\psi^{(N_{\rm nrg},\rm G)}_{\mathcal{Q}}({\mathcal{S}},{\mathcal{S}})}||\rho^{(A)}_{\psi^{(N_{\rm nrg},\rm G)}_{\mathcal{Q}}({\mathcal{S}},-{\mathcal{S}})}\right)(\equiv D_{{\mathcal{S}}{\mathcal{Q}}}(N_{\rm nrg}))$. For ${\mathcal{S}}\neq 0$, the corresponding appropriate ${\mathcal{Q}}$ retained in our NRG calculation are $({\mathcal{S}},{\mathcal{Q}})=(n/2,\pm 1)$, $(n/2,\pm 3)\ (n=1,3,5)$, $(1/2,\pm 5)$,$(3/2,\pm 5)$, $(m,0),(m,\pm 2)$, $(m,\pm 4)\ (m=1,2)$, $(3,0)$. For each $({\mathcal{S}},{\mathcal{Q}})$ mentioned above, the $N_{\rm nrg}$-dependence of relative entropy $D_{{\mathcal{S}}{\mathcal{Q}}}(N_{\rm nrg})$ are shown in Fig.14. The $N_{\rm nrg}$-dependence of $S_{\rm imp}$ is also shown in the figure again as reference. Here, we have confirmed the following inequalities and equality hold for $N$ in the Kondo screening region; $D_{5/2,\pm 3}(N)>D_{2,\pm 4}(N)$$>D_{3,0}(N)>D_{3/2,\pm 5}(N)>D_{5/2,\pm 1}(N)$$>D_{2,\pm 2}(N)>D_{3/2,\pm 3}(N)$ $>D_{2,0}(N)>D_{3/2,\pm 1}(N)$ $>D_{1,\pm 4}(N)>D_{1/2.\pm 5}(N)$ $>D_{1,\pm 2}(N)>D_{1,0}(N)$ $>D_{1/2,\pm 3}(N)>D_{1/2,\pm 1}(N)$(double sign in same order), $D_{SQ}(N)>D_{SQ^{\prime}}(N)\ (|Q|>|Q^{\prime}|)$, and $D_{S,Q}(N)=D_{S,-Q}(N)$. The relationship between the quantum number $({\mathcal{S},Q})$ and the order of the $D_{\mathcal{S},Q}$ values is summarized in Table 3 for ease of reading.
@@
11th ${\mathcal{Q}}=\pm 5$
12th ${\mathcal{Q}}=\pm 2$
13th ${\mathcal{Q}}=0$
@@
It can be seen from Fig.14 that $D_{{\mathcal{S}}{\mathcal{Q}}}(N)$ decreases from $N\simeq 10$ to low temperatures with two continuously connected power laws. Here we define $N_{c}({\mathcal{S},Q})$, the point of maximum curvature of the graph curve of $D_{{\mathcal{S}}{\mathcal{Q}}}(N)$, as the change-point where the power law changes. We find that $N_{c}(\frac{1}{2},\pm 1)=42$ and $N_{c}(\frac{5}{2},\pm 3)=50$, which give the minimum and maximum change-points in $({\mathcal{S},Q})$ considered here, respectively. The positions of those two points are indicated in Fig.14 by arrows. It can been seen that these two points are located within the Kondo screening region indicated by $S_{\rm imp}$. Fig.14 also shows that the change-point for small ${\mathcal{S}}$ is on the high temperature side, while the change-point for larger ${\mathcal{S}}$ is on the low temperature side, mostly.
We next consider the interpretation of the above results. In regions of $N_{\rm nrg}$ where $\langle n\rangle$ and $\langle n_{\uparrow}n_{\downarrow}\rangle$ are not $N_{\rm nrg}$-dependent, the $N_{\rm nrg}$-dependence of $D$ directly reflects the $N_{\rm nrg}$-dependence of $\langle s_{z}\rangle$. In the region of low temperatures below $N_{\rm nrg}\simeq 10$, this is in fact the case, and $D\propto\langle s_{z}\rangle^{2}$. Therefore, the power law of $D$ is a reflection of the power law of $\langle s_{z}\rangle$. The power law of $\langle s_{z}\rangle$ on the high temperature side can be interpreted as the decrease of magnetization due to spin fluctuation in the free spin state, while the power law on the low temperature side can be interpreted as the disappearance of magnetization due to the Kondo effect, i.e. the formation of many-body singlet states. When the value of ${\mathcal{S}}$ is large, the value of $D_{\mathcal{S},Q}$ is large because one would expect $\langle s_{z}\rangle$ to be large, and since states with large ${\mathcal{S}}$ generally correspond to excited states with higher energies, they follow the behavior of lower energy states at lower temperatures, as interpreted by Wilson’s image of renormalization flow. As a result, the change-points for various values of ${\mathcal{S}}$ are widely distributed in the Kondo screening region.
@@ V Conclusion
In order to formulate the distribution of quantum entanglement on quantum impurities and quantum entanglement in quantum impurity pair in the system, for states of interest in quantum impurity systems, we have considered quantum entanglement between a subsystem consisting of one or two quantum impurities arbitrarily selected from the system and its complement, the environmental system. For this purpose, the pure state of interest has been reduced to the subsystem consisting of selected quantum impurities to obtain the density operator, and quantum informative quantities such as entanglement entropy have been calculated. The pure states of interest in quantum impurity systems are often simultaneous eigenstates of physical quantities such as total particle number $Q$, total spin $S_{z}$ ($\Delta Q$), etc. (and often also energy eigenstates of the system). It has been shown that the density operator is diagonal or almost diagonal in several cases. As a result, we have shown that entanglement entropy, mutual information, and relative entropy are given by relatively simple formulas, and that these quantities can be expressed in terms of expectation values of physical quantities on the selected quantum impurities and their correlation functions which can be calculated independently of each other. As we have demonstrated, these expectation values and correlation functions can be calculated with high precision using NRG. It should also be denoted that, apart from their usefulness and possibility of highly accurate numerical calculations, as can be seen within the derivation process, the equations derived here can be applied to systems other than quantum impurity systems(e.g., periodic systems , etc.).
@@
Numerical evaluation of these formulas has been performed for quantum impurity systems consisting of several sites ($2-6$ sites) and SIAM . In the two-site system, we have considered the case of three type of two-body interactions. The results have shown that the maximum quantum entanglement cannot be achieved when the interaction is introduced alone, but if Eq.(79) is satisfied, the three type of two-body interactions can be introduced while maintaining maximum quantum entanglement. Although Eq.(79) itself would be applicable only to two-site systems, it suggests that in systems where several types of two-body interactions exist, the entanglement entropy varies diversely due to competition or cooperation among them. The result is promising for future studies in systems where several types of two-body interactions exist. In the $3-6$ site Hubbard chain, the relationship between its ground state and the entanglement entropy distribution has been studied. In these systems, the entanglement entropy distributions are not flat due to the finite systems with edges. It would be interesting to see how these distributions changes when electron reservoirs are connected to these systems. The behavior of the mutual information in the $4-6$ site Hubbard chains, which is not a monotonically decreasing inter-site distance, also suggests that other systems deserve further study.

### v3

@@ Abstract
We consider quantum entanglement in strongly correlated quantum impurity systems for states manifesting interesting properties such as multi-level Kondo effect and dual nature between itineracy and localization etc.. For this purpose, we set up a system consisting of one or two quantum impurities arbitrarily selected from the system as a subsystem, and investigate quantum entanglement with its environmental system. We reduce the pure state of interest as described above to the subsystem, and formulate quantum informative quantities such as entanglement entropy, mutual information and relative entropy. To demonstrate the potential of the method proposed here, we apply them to a dimer with several types of internal interaction, short Hubbard chains and the single impurity Anderson model to study the relationship between their states and the behaviors of quantum informative quantities. The obtained results suggest that the method proposed here is promising for elucidating the quantum entanglement of pure states in various quantum impurity systems.
@@ I Introduction
Therefore, in this paper, quantum impurities are modeled by Anderson impurities, and in section II, we consider the density operator describing the pure state of interest reduced to the subsystem consisting of one or two selected Anderson impurities in the system. We then formulate quantum informative quantities such as entanglement entropy, mutual information, and relative entropy. In sections III and IV, we apply the formulations in the previous section to a dimer with several types of internal interaction, short Hubbard chains and the Single Impurity Anderson Model(SIAM) to study the relationship between their states and the behaviors of quantum informative quantities. The discussion and conclusions are provided in the last section V.
@@ II.2 Subsystem $A$ consisting of two selected quantum impurities
III Results of application to a dimer with several types of internal interaction, short Hubbard chains
III.1 Dimer with several types of internal interaction
@@
III.2 Short Hubbard chains
@@
The set of the quantum numbers $({\mathcal{S}}_{z},{\mathcal{S}},{\mathcal{Q}})$ of the ground state $\psi^{({\rm G})}_{{\mathcal{S}}_{z},{\mathcal{S}},{\mathcal{Q}}}$ is $({\mathcal{S}}_{z},{\mathcal{S}},{\mathcal{Q}})=\left(0,0,0\right)$ if $M$ is even and $({\mathcal{S}}_{z},{\mathcal{S}},{\mathcal{Q}})=\left(\pm\frac{1}{2},\frac{1}{2},0\right)$ if $M$ is odd. This is because in all cases there are antiferromagnetic correlations in the ground state, but the magnetic moments cancel each other when $M$ is even, and there is an unpaired magnetic moment when $M$ is odd.
@@ IV.2 Quantum entanglement in energy eigenstates depending on $N_{\rm nrg}$
We next consider two states with ${\mathcal{S}}_{z}={\mathcal{S}}$ and ${\mathcal{S}}_{z}=-{\mathcal{S}}$ for the lowest energy state $\psi^{(N_{\rm nrg},\rm G)}_{\mathcal{Q}}({\mathcal{S}},{\mathcal{S}}_{z})$ within ${\mathcal{S}}\neq 0$ and the corresponding appropriate ${\mathcal{Q}}$ for each $N_{\rm nrg}$, and then reduce these two states to the Anderson impurity system and consider the relative entropy $D\left(\rho^{(A)}_{\psi^{(N_{\rm nrg},\rm G)}_{\mathcal{Q}}({\mathcal{S}},{\mathcal{S}})}||\rho^{(A)}_{\psi^{(N_{\rm nrg},\rm G)}_{\mathcal{Q}}({\mathcal{S}},-{\mathcal{S}})}\right)(\equiv D_{{\mathcal{S}}{\mathcal{Q}}}(N_{\rm nrg}))$. For ${\mathcal{S}}\neq 0$, the corresponding appropriate ${\mathcal{Q}}$ retained in our NRG calculation are $({\mathcal{S}},{\mathcal{Q}})=(n/2,\pm 1)$, $(n/2,\pm 3)\ (n=1,3,5)$, $(1/2,\pm 5)$,$(3/2,\pm 5)$, $(m,0),(m,\pm 2)$, $(m,\pm 4)\ (m=1,2)$, $(3,0)$. For each $({\mathcal{S}},{\mathcal{Q}})$ mentioned above, the $N_{\rm nrg}$-dependence of relative entropy $D_{{\mathcal{S}}{\mathcal{Q}}}(N_{\rm nrg})$ are shown in Fig.14. The $N_{\rm nrg}$-dependence of $S_{\rm imp}$ is also shown in the figure again as reference. Here, we have confirmed the following inequalities and equality hold for $N$ in the Kondo screening region; $D_{5/2,\pm 3}(N)>D_{3,0}(N)$$>D_{2,\pm 4}(N)>D_{3/2,\pm 5}(N)>D_{5/2,\pm 1}(N)$$>D_{2,\pm 2}(N)>D_{2,0}(N)$ $>D_{3/2,\pm 3}(N)>D_{3/2,\pm 1}(N)$ $>D_{1,\pm 4}(N)>D_{1.\pm 2}(N)$ $>D_{1,0}(N)>D_{1/2,\pm 5}(N)$ $>D_{1/2,\pm 3}(N)>D_{1/2,\pm 1}(N)$ (double sign in same order), $D_{SQ}(N)>D_{SQ^{\prime}}(N)\ (|Q|>|Q^{\prime}|)$, and $D_{S,Q}(N)=D_{S,-Q}(N)$.
The relationship between the quantum number $({\mathcal{S},Q})$ and the order of the $D_{\mathcal{S},Q}$ values is summarized in Table 3 for ease of reading.
@@
11th ${\mathcal{Q}=\pm 2}$
12th ${\mathcal{Q}}=0$
13th ${\mathcal{Q}}=\pm 5$
@@
It can be seen from Fig.14 that $D_{{\mathcal{S}}{\mathcal{Q}}}(N)$ decreases from $N\simeq 10$ to low temperatures with two continuously connected power laws. Here we define $N_{c}({\mathcal{S},Q})$, the point of maximum curvature of the graph curve of $D_{{\mathcal{S}}{\mathcal{Q}}}(N)$, as the change-point where the power law changes. We find that $N_{c}(\frac{1}{2},\pm 1)=44$ and $N_{c}(\frac{5}{2},\pm 3)=50$, which give the minimum and maximum change-points in $({\mathcal{S},Q})$ considered here, respectively. The positions of those two points are indicated in Fig.14 by arrows. It can been seen that these two points are located within the Kondo screening region indicated by $S_{\rm imp}$. Fig.14 also shows that the change-point for small ${\mathcal{S}}$ is on the high temperature side, while the change-point for larger ${\mathcal{S}}$ is on the low temperature side, mostly.
We next consider the interpretation of the above results. In regions of $N_{\rm nrg}$ where $\langle n\rangle$ and $\langle n_{\uparrow}n_{\downarrow}\rangle$ are not $N_{\rm nrg}$-dependent, the $N_{\rm nrg}$-dependence of $D$ directly reflects the $N_{\rm nrg}$-dependence of $\langle s_{z}\rangle$. In the region of low temperatures below $N_{\rm nrg}\simeq 10$, this is in fact the case, and $D\propto\langle s_{z}\rangle^{2}$. Therefore, the power law of $D$ is a reflection of the power law of $\langle s_{z}\rangle$. The power law of $\langle s_{z}\rangle$ on the high temperature side due to scale invariance of the free spin regime shows the decrease of magnetization due to spin fluctuation in the free spin state, while the power law on the low temperature side due to scale invariance of the Kondo regime shows the disappearance of magnetization due to the Kondo effect, i.e. the formation of many-body singlet states. When the value of ${\mathcal{S}}$ is large, the value of $D_{\mathcal{S},Q}$ is large because one would expect $\langle s_{z}\rangle$ to be large, and since states with large ${\mathcal{S}}$ generally correspond to excited states with higher energies, they follow the behavior of lower energy states at lower temperatures, as interpreted by Wilson’s image of renormalization flow. As a result, the change-points for various values of ${\mathcal{S}}$ are widely distributed in the Kondo screening region.
@@ V Conclusion
In order to formulate the distribution of quantum entanglement on quantum impurities and quantum entanglement in quantum impurity pair in the system, for states of interest in quantum impurity systems, we have considered quantum entanglement between a subsystem consisting of one or two quantum impurities arbitrarily selected from the system and its complement, the environmental system. For this purpose, the pure state of interest has been reduced to the subsystem consisting of selected quantum impurities to obtain the density operator, and quantum informative quantities such as entanglement entropy have been calculated. The pure states of interest in quantum impurity systems are often simultaneous eigenstates of physical quantities such as total particle number $Q$, total spin $S_{z}$ ($\Delta Q$), etc. (and often also energy eigenstates of the system). It has been shown that the density operator is diagonal or almost diagonal in several cases. As a result, we have shown that entanglement entropy, mutual information, and relative entropy are given by relatively simple formulas, and that these quantities can be expressed in terms of expectation values of physical quantities on the selected quantum impurities and their correlation functions which can be calculated independently of each other. As we have demonstrated, these expectation values and correlation functions can be calculated with high precision using NRG. It should also be denoted that, apart from their usefulness and possibility of highly accurate numerical calculations, as can be seen within the derivation process, the equations derived here can be applied to systems other than those presented in this paper(e.g., periodic systems , etc.).
@@
Numerical evaluation of these formulas has been performed for the dimer with several types of internal interaction, short Hubbard chains and SIAM . In the dimer with several types of internal interaction, we have considered the case of three type of two-body interactions. The results have shown that the maximum quantum entanglement cannot be achieved when the interaction is introduced alone, but if Eq.(79) is satisfied, the three type of two-body interactions can be introduced while maintaining maximum quantum entanglement. Although Eq.(79) itself would be applicable only to dimers, it suggests that in systems where several types of two-body interactions exist, the entanglement entropy varies diversely due to competition or cooperation among them. The result is promising for future studies in systems where several types of two-body interactions exist. In the $3-6$ site Hubbard chain, the relationship between its ground state and the entanglement entropy distribution has been studied. In these systems, the entanglement entropy distributions are not flat due to the finite systems with edges. It would be interesting to see how these distributions changes when electron reservoirs are connected to these systems. The behavior of the mutual information in the $4-6$ site Hubbard chains, which is not a monotonically decreasing inter-site distance, also suggests that other systems deserve further study.